### v2

# Unveiling the Hidden: Movie Genre and User Bias in Spoiler Detection

Abstract.
Spoilers in movie reviews are important on platforms like IMDb and Rotten Tomatoes, offering benefits and drawbacks. They can guide some viewers’ choices but also affect those who prefer no plot details in advance, making effective spoiler detection essential. Existing spoiler detection methods mainly analyze review text, often overlooking the impact of movie genres and user bias, limiting their effectiveness. To address this, we analyze movie review data, finding genre-specific variations in spoiler rates and identifying that certain users are more likely to post spoilers. Based on these findings, we introduce a new spoiler detection framework called GUSD  [^1] The code is available at https://github.com/AI-explorer-123/GUSD (Genre-aware and User-specific Spoiler Detection), which incorporates genre-specific data and user behavior bias. User bias is calculated through dynamic graph modeling of review history. Additionally, the R2GFormer module combines RetGAT (Retentive Graph Attention Network) for graph information and GenreFormer for genre-specific aggregation. The GMoE (Genre-Aware Mixture of Experts) model further assigns reviews to specialized experts based on genre. Extensive testing on benchmark datasets shows that GUSD achieves state-of-the-art results. This approach advances spoiler detection by addressing genre and user-specific patterns, enhancing user experience on movie review platforms.
Spoiler Detection, Movie Genre, User Bias, Mixture-of-Experts
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
1. Introduction
Spoilers in movie reviews have become an important component of the movie-viewing experience on popular platforms like IMDb and Rotten Tomatoes (Cao et al., 2019). For those who hope to learn the plot of the movie in advance to judge whether they like it or not, the spoilers are helping, while for those who prefer to experience a movie without prior knowledge, the spoilers can severely diminish the enjoyment by revealing crucial plot points, undermining suspense, and eliciting negative emotions among viewers (Loewenstein, 1994). Thus, effective spoiler detection methods are crucial for maintaining a positive user experience.
Figure 1. An illustrative example of the data used in our spoiler detection study. The image shows a review of the movie Moana, including the movie’s genres (Adventure, Animation), summary, the review’s content, and user-specific details. All reviews from the user are color-coded: blue indicates non-spoiler content, while red indicates spoiler content.
Existing spoiler detection methods primarily focus on the textual content of reviews. For example, DNSD (Chang et al., 2018) integrates review sentences and movie genres, while SpoilerNet (Wan et al., 2019) utilizes a Hierarchical Attention Network and incorporates the item-specificity information. More recent approaches, such as MVSD (Wang et al., 2023), incorporate advanced techniques like syntax-aware graph neural networks and external movie knowledge to improve detection performance. Nevertheless, these methods still exhibit notable limitations. Solely relying on textual content proves insufficient for robust spoiler detection (Wang et al., 2023). Moreover, spoilers are often genre-specific, with varying characteristics depending on the movie’s genre — for instance, suspense films focus on plot details, whereas action films emphasize fight scenes. As a result, two significant challenges in spoiler detection remain unaddressed:
  • Diverse Genres. Previous works have largely ignored the impact of movie genres on the spoiler rate. Our analysis of the dataset indicates substantial differences in spoiler rate across genres, with specific categories defined according to IMDb standards. As shown in Figure 2(a), movies that heavily rely on plot twists and suspense, such as Film-Noir and Adventure, are more prone to having spoilers in reviews compared to genres like Musical or Documentary. This is understandable since plot-driven movies tend to have more critical plot points that can be spoiled. This variation in spoiler rate demonstrates the importance of considering genre-specific characteristics when developing spoiler detection models. By incorporating genre information, we could better capture these differences and improve the performance of spoiler detection.
  • User-specific Behavior Bias. User behavior varies significantly, with some users being more prone to posting spoilers than others. Our statistical analysis shows a clear trend where certain users tend to post spoiler reviews more frequently. As illustrated in Figure 2(b), the graph of user spoiler rate distribution shows that a large proportion of users post very few spoilers, while a smaller, yet significant, group of users frequently post spoilers. This distribution indicates a noticeable user bias, highlighting that certain users are more likely to post spoilers than others. Leveraging these user-specific behavior bias can improve the detection performance by allowing models to adapt to user behavior bias.
Figure 2. (a) Spoiler rate across different movie genres (partial) in LCS dataset. (b) Kernel density estimation plot and distribution histogram of spoiler across different users.
To address the challenges of genre-specific spoiler tendencies and user bias in spoiler detection, we propose a comprehensive framework named GUSD (Genre-aware and User-specific Spoiler Detection). This framework integrates genre information, user behavior bias, and global perception GNN. Our method begins with preprocessing movie, user, and review data. Then user bias is captured from review history through dynamic graph modeling. After that, the core component, R2GFormer (RetGAT and GenreFormer), which consists of RetGAT (Retentive Graph Attention Network) and GenreFormer, processes the graph information. RetGAT aggregates all the features globally, and GenreFormer enhances representation by integrating genre features, allowing for comprehensive genre-specific and cross-genre interactions. And an Aggregator combines these features, and then the GMoE (Genre-Aware Mixture of Experts) model assigns reviews to different experts based on their corresponding movie genres, improving traditional MoE model performance. Finally, a classifier performs spoiler detection using the aggregated features.
Extensive experiments show that GUSD achieves state-of-the-art performance. We also conduct robustness studies, ablation studies, and specific experiments on GMoE and user bias to validate our proposed modules’ effectiveness.
Our main contributions are summarized as follows:
  • We are the first to model the complex interactions between genre-specific information and long-term user review behaviors for spoiler detection, providing a nuanced approach by understanding genre-specific spoiler characteristics and leveraging user behavior bias.
  • We propose the GUSD framework, a novel spoiler detection system that integrates several key components: the GenreFormer to capture genre-specific spoiler tendencies, the GMoE model to dynamically assign reviews based on genres, and dynamic graph modeling to capture user bias. This cohesive integration enhances overall accuracy and robustness.
  • Our method GUSD achieves state-of-the-art performance in spoiler detection. Extensive experiments on two benchmark datasets demonstrate its robustness and effectiveness, showing superior performance across various conditions.
Figure 3. Overview of our proposed GUSD framework, which integrates genre-specific information, user behavior bias, and global perceptive RetGAT for spoiler detection. It preprocesses movie, user, and review data with MLP and language models, and captures user bias via dynamic graph modeling. Then the data is processed by the R2GFormer component. An Aggregator merges these features, and then GMoE assigns reviews to experts based on genres. Finally, a classifier detects spoilers using the aggregated features.
2. Related Work
2.1. Spoiler Detection
The goal of automatic spoiler detection is to identify spoilers in reviews from domains like television (Boyd-Graber et al., 2013), books (Wan et al., 2019), and movies (Boyd-Graber et al., 2013). Existing approaches to spoiler detection can be broadly classified into three categories: keyword matching methods, machine learning techniques, and deep learning models.
2.1.1. Keyword matching methods.
These approaches rely on a set of predefined keywords to identify spoilers. Examples include keywords related to sports teams or events (Nakamura and Tanaka, 2007), or actors’ names (Golbeck, 2012). Although useful in specific scenarios, this method requires manual keyword definition and lacks generalizability across different application contexts.
2.1.2. Machine learning techniques.
These methods often involve topic modeling or support vector machines using handcrafted features. For example, Guo et al. (Guo and Ramakrishnan, 2010) applied a bag-of-words representation combined with an LDA-based model for spoiler detection. Jeon et al. (Jeon et al., 2013) developed an SVM classifier incorporating four extracted features, while Boyd et al. (Boyd-Graber et al., 2013) utilized lexical features and meta-data of review subjects (e.g., movies and books) in an SVM model.
2.1.3. Deep learning models.
These models mainly leverage NLP techniques, employing RNNs, LSTMs, Transformer, and language models to process review texts and movie information through end-to-end training. Bao et al. (Bao et al., 2021) utilized LSTMs, BERT, and RoBERTa for sentence-level spoiler detection. DNSD (Chang et al., 2018) focused on incorporating external genre information using GRU and CNN. SpoilerNet (Wan et al., 2019) introduced item-specificity and bias with bi-RNN enhanced by GRU. SDGNN (Chang et al., 2021) leveraged dependency relations between context words in sentences with graph neural networks to capture semantics.
While some existing methods incorporate genre information and user bias  (Chang et al., 2018; Wróblewska et al., 2021), they often rely on quite simple techniques such as concatenating or adding these additional features to the initial review features. Such approaches lack the sophistication needed for more effective and intricate modeling of genre features and user biases.
2.2. Mixture of Experts
The Mixture of Experts (MoE) approach, grounded in the Divide-and-Conquer principle, segments an input sample into sub-tasks and trains specialized experts for each sub-task. This method is extensively utilized in NLP to boost model capacity (Shazeer et al., 2017) and enhance reasoning capabilities (Madaan et al., 2021). Shazeer et al. (Shazeer et al., 2017) introduced a sparsely-gated Mixture-of-Experts layer, enabling conditional computing in large language models. Fedus et al. (Fedus et al., 2022) developed simplified routing algorithms for MoE to enhance training stability and reduce computational costs. Furthermore, Soft-MoE (Puigcerver et al., 2023) was introduced to mitigate issues like training instability and token dropping inherent in traditional MoE approaches.
Despite these advancements, traditional MoE methods assign tokens dynamically, which can cause incorrect token assignment, particularly when the dataset contains explicit category information such as movie genres and their associated reviews.
3. Methodology
Figure 3 shows the architecture of our proposed GUSD framework. This framework integrates genre-specific information, user behavior bias, and global receptive RetGAT for spoiler detection. Specifically, movie, user, and review data are first to be preprocessed, while user bias is extracted from review history using dynamic graph modeling. The R2GFormer component, consisting of RetGAT and GenreFormer, then processes graph features. RetGAT aggregates graph data, while GenreFormer handles genre-specific data. An Aggregator aggregates these features. The GMoE model assigns reviews to experts based on their related movie genres. Finally, a classifier utilizes the aggregated features to detect spoilers.
3.1. Data preprocessing
3.1.1. Meta information.
For review nodes, user nodes, and movie nodes, each type possesses meta information. After collecting the meta information for all three types of nodes, we pad them to the same length. A two-layer MLP is then employed as the meta encoder, producing the meta embeddings $N_{m}$.
3.1.2. Textual information.
The textual content is fundamental for effective spoiler detection. To generate high-quality embeddings, we leverage an LM as our text encoder. Specifically, we augment the initial textual content with the textual descriptions of the node’s metadata. This augmentation enriches the embeddings by providing additional contextual information about the node. Subsequently, we employ the LM to encode the nodes’ textual information. The encoded embeddings are then transformed using a single-layer MLP, producing refined embeddings $N_{t}$.
3.1.3. User bias acquisition using dynamic-graph pre-training.
To better capture the dynamic information attributes of users, we adopt the dynamic graph to handle users’ review history flexibly. Thus, we need to convert the static dataset into a dynamic format. We utilize the given connections between different nodes and the time information of the reviews to form the dynamic event stream. The details about the formation of the graph will be displayed in Section 3.2. Considering the absence of additional information about the dynamic edges, we simply initial the edge features as zero vectors. Then, we employ the robust DyGFormer (Yu et al., 2023) as our dynamic graph encoder to capture the interactions among various nodes and utilize Link Prediction as the downstream task. We specifically obtain the features of user nodes as user bias $U_{b}$ from the dynamic graph encoder.
3.2. R2GFormer
After acquiring the initial meta and textual features of the nodes, we use the textual features $N_{t}$, which include rich information about the nodes, as the initial embeddings $N_{g}^{(0)}$ for R2GFormer. This graph encoder not only models the complex relations and interactions between users, reviews, and movies but also handles features from different genres of the entire graph and captures remote dependencies. First, we will introduce how we construct the whole graph. Then, we decompose an R2GFormer layer into two parts: RetGAT and Genreformer, which will be introduced respectively.
3.2.1. Graph Construction.
We first construct a directed graph consisting of three types of nodes: { User, Review, Movie } and the following three types of edges:
E1: Movie-Review We connect a review node to a movie node if the review is about the movie, but not vice versa. This setup allows movie information to influence the review while ensuring that the review information does not affect the movie.
E2: Review-User We connect a review node to a user node if the review is posted by the user.
E3: User-Review We connect a user node to a review node if the user posts the review.
3.2.2. RetGAT.
Inspired by the work of (Sun et al., 2023; Fan et al., 2024; Nikolentzos et al., 2020), we propose RetGAT, which extends the RetNet framework by integrating a global perception into GAT, incorporating explicit exponential decay for nodes within $k$ hops and truncation for nodes beyond $k$ hops. This method ensures a broad receptive field while balancing performance and computational complexity by dynamically adjusting the influence of nodes based on their distance, with closer nodes having a higher impact on the final node features.
To achieve this global receptive field, we utilize $k$ parallel GAT layers to separately aggregate information from $k$-hop neighbors. For each layer, node features are aggregated from different $k$-hop neighborhoods, applying decay factors to control the influence of information from various hops. Beyond $k$ hops, the influence of nodes is truncated to maintain computational efficiency and focus on relevant information within the $k$-hop range. Note that we correct the algorithm that previous work (Atwood and Towsley, 2016; Wang et al., 2020) used to compute $k$-hop neighbors, and the details are available in the supplementary material.
The decay factor $\delta_{h}$ for hop $h$ ($h\leq k$) is defined as:
(1) $\delta_{h}$ $=\exp(-\alpha\cdot h),$
where $\alpha$ is a hyperparameter controlling the intensity of exponential decay. For the $l$-th layer, the contribution of the $h$-th hop neighbors is computed as:
[…]
where $\mathcal{A}_{h}$ represents the adjacency matrix at the $h$-th hop, and ${N_{g}}^{(l-1)}$ denotes the input node features of the $(l-1)$-th layer.
The final node features ${N_{g}}^{(l)}$ for layer $l$ are aggregated from all the $k$-hop contributions using an Aggregator function:
(3) ${N_{g}^{(l)}}$ $=\text{AGGREGATOR}_{r}^{(l)}(\text{${N_{g,1}^{(l)}}$},\text{${N_{g,2}^{(l)}}$},\ldots,\text{${N_{g,h}^{(l)}}$},\ldots,\text{${N_{g,k}^{(l)}}$}).$
where $\text{AGGREGATOR}_{r}^{(l)}$ is the aggregator for RetGAT at the l-th layer, which can be summation, concatenation, or a TransformerEncoder (TRM).
3.2.3. Genreformer.
After message passing among different $k$-hop neighbors, Genreformer aims to pay attention to the entire graph to extract more comprehensive features of different genres. First, the global genre-specific representation is obtained by aggregating the features of all the review and movie nodes belonging to the same genre in the $l$-th layer, which is as follows:
(4) $g_{j}^{(l)}=\text{AGGREGATOR}^{(l)}_{g}(\{n_{g,i}^{(l)}\mid i\in\mathcal{N}_{j}\}),$
where $n_{g,i}^{(l)}$ is the feature of node $i$ in the $l$-th layer, $\mathcal{N}_{j}$ denotes the set of nodes (both review and movie nodes) belonging to the $j$-th genre, and $\text{AGGREGATOR}^{(l)}_{g}$ is the aggregator of Genreformer in the $l$-th layer, which can be summation, concatenation, or a TRM.
Next, we use a TRM to facilitate inter-genre information interaction, allowing each genre to acquire information from similar genres to further enrich its features:
(5) $[g_{1}^{(l)}g_{2}^{(l)}\cdots g_{j}^{(l)}\cdots g_{c}^{(l)}]=\text{TRM}([g_{1}^{(l)}g_{2}^{(l)}\cdots g_{j}^{(l)}\cdots g_{c}^{(l)}]),$
where $c$ is the number of genres.
After interaction among genres, the genre features are fused with the movies and reviews nodes features. Then we incorporate the nodes with their genre features. Since some movies and reviews cover more than one genre, we take the average of the genre features a node covers. Subsequently, we concatenate the review or movie feature and its genre feature, then use an MLP to project it into the desired feature space, $i.e.$,
(6) $z_{i}^{(l)}$ $=\text{MEAN}(\{g_{j}^{(l)}\mid j\in G_{i}\}),$
(7) $n_{g,i}^{(l)}$ $=\text{\text{MLP}}([n_{g,i}^{(l)}\parallel z_{i}^{(l)}]).$
where $G_{i}$ denotes the set of genres to which node $i$ belongs, $n_{g,i}^{(l)}$ denotes the feature of node $i$ in the $l$-th layer, $z_{i}^{(l)}$ is the aggregated genre feature for node $i$ in the $l$-th layer, and $\parallel$ means concatenation. The MLP projects the concatenated feature into the desired feature space.
3.2.4. Overall interaction.
One layer of our proposed R2GFormer layer, however, cannot enable the information interaction between all information sources. In order to further facilitate the interaction among the nodes, we employ $L\times$ R2GFormer layers for node representation learning. The representation of the nodes is updated after each layer, incorporating information from different sources. This process can be formulated as follows:
(8) $N_{g}^{(l)}=\text{R2GFormer}(N_{g}^{(l-1)},\mathcal{E},G).$
where $\mathcal{E}$ denotes all the edges of the sampling graph, $G$ denotes the genres of every movie and review node. After $L\times$ R2GFormer layers, we obtain the final graph representation $N_{g}^{(L)}$ of all nodes.
3.3. Multimodal Fusion
Through the data above processed by the R2GFormer, we have obtained the graph structural information and meta information for all nodes, including user, review and movie nodes. The next step is to fuse the multimodal information.
For each type of node, we utilize a type-wise TRM to facilitate inter-modal information interaction, then concatenate features of different modals followed by an MLP to get the final representation for each node, i.e.,
(9) $U_{g},U_{m},U_{b}=\text{TRM}([U_{g}\ U_{m}\ U_{b}]$ $,U=\text{\text{MLP}}([U_{g}\parallel U_{m}\parallel U_{b}])),$
(10) $R_{g},R_{m}=\text{TRM}([R_{g}\ R_{m}]$ $,R=\text{\text{MLP}}([R_{g}\parallel R_{m}])),$
(11) $M_{g},M_{m}=\text{TRM}([M_{g}\ M_{m}]$ $,M=\text{\text{MLP}}([M_{g}\parallel M_{m}]),$
where $U_{g}$, $R_{g}$ and $M_{g}$ are derived from $N_{g}$, $U_{m}$, $R_{m}$ and $M_{m}$ are dervied from $N_{m}$.
After obtaining the comprehensive representation of each type of node, we then concatenate each review feature $r_{i}$ with its corresponding movie feature $m_{i}$ and user feature $u_{i}$:
(12) $r_{i}=[\ r_{i}\parallel m_{i}\parallel u_{i}\ ].$
3.4. GMoE
Inspired by the successful applications of Mixture-of-Experts in NLP and bot detection, and its capability to handle the small subsets of the whole dataset, we adopt MoE to handle different genres of reviews. However, distinct from the latent subsets of the dataset in the classic MoE application scenario, our datasets already have the genre information of movies, as well as their related reviews. So we improve MoE to the proposed GMoE.
Specifically, instead of using the gating mechanism in the traditional MoE structure, we assign tokens to experts simply according to their genres: which genre it belongs to, which expert will deal with it; how many genres it belongs to, how many experts will deal with it.
(13) $r_{i}=\text{AGGREGATOR}_{m}(\{\text{Expert}_{j}(r_{i})\mid\forall{j}\in{G_{i}}\}).$
where $G_{i}$ denotes the set of genres to which node $i$ belongs; AGGRE-
GATORm can be summation, concatenation, or a TRM; each Expert is a MLP for simplicity.
3.5. Learning and Optimization
After using GMoE to process genre-specific information, we acquire the final representation $r_{i}$ for the $i$-th review. Then we apply a linear transformation to $r_{i}$ to obtain spoiler detection result $\hat{y_{i}}$. To train GUSD, We optimize the network by cross-entropy loss with $L_{2}$ regularization. The total loss function is as follows:
(14) $Loss=-\sum_{i\in\mathcal{R}}{y_{i}\log\hat{y_{i}}}+\lambda\sum_{\theta\in\Theta}{\theta^{2}}.$
where $\hat{y_{i}}$ and $y_{i}$ are the prediction for the $i$-th review and its corresponding ground truth, respectively. $\mathcal{R}$ encompasses all the reviews in the training set, while $\Theta$ denotes all trainable model parameters in GUSD, and $\lambda$ is a hyper-parameter that maintains the balance between the two parts.
Table 1. Accuracy, AUC, and binary F1-score of GUSD and three types of baseline methods on two spoiler detection datasets. We run all experiments five times to ensure a consistent evaluation and report the average performance as well as standard deviation in parentheses. Bold indicates the best performance, underline the second best. GUSD consistently outperforms the three types of methods on both benchmarks.
Model Kaggle LCS
F1 AUC Acc F1 AUC Acc
BERT 44.02 $\pm$1.09 63.46 $\pm$0.46 77.78 $\pm$0.09 46.14 $\pm$2.84 64.82 $\pm$1.36 79.96 $\pm$0.38
RoBERTa 50.93 $\pm$0.76 66.94 $\pm$0.40 79.12 $\pm$0.10 47.72 $\pm$0.44 65.55 $\pm$0.22 80.16 $\pm$0.03
BART 46.89 $\pm$1.55 64.88 $\pm$0.71 78.47 $\pm$0.09 48.18 $\pm$1.22 65.79 $\pm$0.62 80.14 $\pm$0.07
DeBERTa 49.94 $\pm$1.13 66.42 $\pm$0.59 79.08 $\pm$0.09 47.38 $\pm$2.22 65.42 $\pm$1.08 80.13 $\pm$0.08
Bge-Large 52.51 $\pm$0.58 67.74 $\pm$0.37 77.44 $\pm$0.21 52.68 $\pm$0.36 68.46 $\pm$0.23 79.24 $\pm$0.11
GCN 59.22 $\pm$1.18 71.61 $\pm$0.74 82.08 $\pm$0.26 62.12 $\pm$1.18 73.72 $\pm$0.89 83.92 $\pm$0.23
R-GCN 63.07 $\pm$0.81 74.09 $\pm$0.60 82.96 $\pm$0.16 62.99 $\pm$0.89 76.18 $\pm$0.72 85.19 $\pm$0.21
GAT 60.98 $\pm$0.09 72.72 $\pm$0.60 82.43 $\pm$0.01 65.73 $\pm$0.12 75.92 $\pm$0.13 85.18 $\pm$0.02
SimpleHGN 60.12 $\pm$1.04 71.60 $\pm$0.88 82.08 $\pm$0.26 63.79 $\pm$0.88 74.64 $\pm$0.64 84.66 $\pm$1.61
HGT 63.99 $\pm$0.25 75.61 $\pm$0.25 81.66 $\pm$0.23 60.89 $\pm$0.46 73.96 $\pm$0.53 81.86 $\pm$0.16
GPS 61.04 $\pm$0.84 73.50 $\pm$0.53 81.25 $\pm$0.55 64.21 $\pm$0.30 75.60 $\pm$0.92 82.40 $\pm$0.91
DNSD 46.33 $\pm$2.37 64.50 $\pm$1.11 78.44 $\pm$0.14 44.69 $\pm$1.64 64.10 $\pm$0.74 79.76 $\pm$0.08
SpoilerNet 57.19 $\pm$0.69 70.64 $\pm$0.44 79.85 $\pm$0.10 62.86 $\pm$0.38 74.62 $\pm$0.69 83.23 $\pm$0.23
MVSD 65.08 $\pm$0.69 75.42 $\pm$0.56 83.59 $\pm$0.11 69.22 $\pm$0.61 78.26 $\pm$0.63 86.37 $\pm$0.08
Ours 80.24 $\pm$0.73 87.00 $\pm$0.37 89.65 $\pm$0.36 75.37 $\pm$0.10 83.71 $\pm$0.27 88.32 $\pm$0.08
4. Experiment
4.1. Experiment Settings
4.1.1. Dataset.
To evaluate our GUSD framework along with 14 other representative baselines on two widely recognized datasets: LCS (Wang et al., 2023) and Kaggle (Wan et al., 2019):
  • LCS is a comprehensive dataset for automatic spoiler detection, comprising 1,860,715 reviews, 259,705 users, and 147,191 movies. And about 24.59% (457,500) of the reviews are spoilers.
  • Kaggle, introduced in 2019, consists of 573,913 valid reviews, 293,407 users, and 1,572 movies. And about 25.87% (150,924) of the reviews are spoilers.
Note that both datasets include the genre information of all movies, with specific categories defined according to IMDb standards, which facilitates the operation of our GUSD framework. Following MVSD (Wang et al., 2023), we randomly split the reviews into training, validation, and test sets with a ratio of 7:2:1.
4.1.2. Baselines.
To achieve a comprehensive evaluation, we compare GUSD with pre-trained language models, GNN-based models, and task-specific baselines. For the pre-trained language models, the procedure involves feeding the review text into the model, averaging all token embeddings, and then applying two fully connected layers to perform spoiler detection. Regarding the GNN-based models, the graph neural network takes the output of RoBERTa (Liu, 2019) as the initial node features. Below, we provide a concise overview of each baseline method.
  • BERT (Devlin, 2018) is a language model pre-trained on extensive natural language data, using masked language modeling and next sentence prediction tasks.
  • RoBERTa (Liu, 2019) improves upon BERT by eliminating the next sentence prediction task and enhancing masking techniques.
  • BART (Lewis, 2019) is a pre-trained language model that advances traditional autoregressive models through bidirectional encoding and denoising objectives.
  • DeBERTa (He et al., 2021) refines BERT by implementing disentangled attention and an improved mask decoder, making it a more advanced language model.
  • Bge-Large (Xiao et al., 2023) is trained on a comprehensive training dataset C-MTP, combining vast unlabeled data and diverse labeled data.
  • GCN (Kipf and Welling, 2016) is a foundational graph neural network that performs convolutions on graph nodes and their neighbors, effectively propagating information.
  • R-GCN (Schlichtkrull et al., 2018) extends GCN to handle multi-relational graphs by incorporating relation-specific weights.
  • GAT (Veličković et al., 2017) is a graph neural network that applies attention mechanisms to dynamically assign importance to neighboring nodes.
  • SimpleHGN (Lv et al., 2021) is tailored for heterogeneous graphs, integrating multiple types of nodes and edges with a shared embedding space and adaptive aggregation strategies.
  • GPS (Rampášek et al., 2022) propose a recipe to build a general, powerful, scalable graph Transformer with linear complexity.
  • HGT (Hu et al., 2020) design node- and edge-type dependent parameters to characterize the heterogeneous attention over each edge for modeling Web-scale heterogeneous graphs.
  • DNSD (Chang et al., 2018) is a spoiler detection method that employs a CNN-based genre-aware attention mechanism.
  • SpoilerNet (Wan et al., 2019) uses a hierarchical attention network and GRU alongside item and user bias terms for spoiler detection.
  • MVSD (Wang et al., 2023) leverages external movie knowledge and user networks to detect spoilers.
4.2. Main Results
We evaluated our GUSD framework and 14 other baselines on two datasets. The results presented in Table 1 demonstrate the following:
  • GUSD consistently outperforms all baselines across both datasets. Specifically, compared with the previous state-of-the-art method MVSD (Wang et al., 2023), GUSD achieves 6.1% higher Binary-F1, 5.5% higher AUC, and 2.0% higher accuracy on the LCS dataset, as well as 15.2% higher Binary-F1, 11.6% higher AUC, and 6.1% higher accuracy on the Kaggle dataset. These improvements are statistically significant.
  • In both datasets, graph-based models generally outperform other types of baselines, reaffirming the importance of analyzing the graph structure of reviews and their corresponding users and movies.
  • Compared with DNSD (Chang et al., 2018), which also focuses on genre features of the reviews, GUSD surpasses DNSD across all three metrics in both datasets, further proving the effectiveness and robustness of our global-aware genreformer and GMoE methods.
  • Both SpoilerNet (Wan et al., 2019) and GUSD utilize user bias, but GUSD outperforms SpoilerNet in all three metrics across both datasets. This demonstrates that our dynamic graph pretraining can better identify the latent behavior pattern of whether a user is likely to post spoilers.
4.3. Ablation Study
As GUSD outperforms all the baselines and has reached state-of-the-art (SOTA) performance across the two datasets, we conducted ablation study to further explore the impact of each part of GUSD on the final performance with the Kaggle Dataset. The results are shown in Table 2.
  • To assess the importance of user bias information, we removed the user bias component $U_{b}$. The results in Table 2 show an obvious decrease in performance, confirming that user bias information is critical for effective spoiler detection.
  • For the RetGAT component, we evaluated two variations: using an approximate method to compute $k$-hop neighbors (Atwood and Towsley, 2016; Wang et al., 2020) and replacing our RetGAT with a standard GAT. Both variations lead to a drop in performance, indicating the necessity of our RetGAT design for capturing appropriate graph structures.
  • To investigate the impact of GMoE, we replace it with a simple MLP, a traditional MoE (Shazeer et al., 2017), and a Soft-MoE (Puigcerver et al., 2023). Note that we set the number of experts of the traditional MoE and Soft-MoE to be the same as in GMoE, i.e., the number of genres. From the results shown in Table 2, the full model with GMoE achieves the best performance, highlighting the effectiveness of our GMoE design in handling explicit genre information. Moreover, replacing the GMoE with an MLP performs the worst, proving the rationality of using genre information.
  • We also evaluate the Genreformer component by removing it and replacing the mean AGGREGATOR with sum pooling and max pooling. The results in Table 2 show that the full Genreformer with sum AGGREGATOR outperforms these ablated versions, validating the necessity of the Genreformer.
Table 2. Ablation study of GUSD on Kaggle Dataset. Bold indicates the best performance, underline the second best.
Catagory Ablation Settings F1 AUC Acc
User bias -w/o $U_{b}$ 78.85 85.78 88.58
RetGAT approximate way 78.40 85.39 88.39
normal GAT 78.08 85.22 88.29
GMoE MLP 78.52 86.09 88.25
MoE (Shazeer et al., 2017) 78.75 86.36 88.54
Soft MoE (Puigcerver et al., 2023) 79.10 86.10 88.78
Genreformer -w/o genreformer 77.73 84.68 88.13
sum pooling 79.84 86.56 89.28
max pooling 78.77 85.52 89.13
Ours GUSD 80.24 87.00 89.65
4.4. GMoE Study
In this section, we conduct further experiments on the GMoE to better understand its effectiveness compared to other previous MoE methods such as traditional MoE (Shazeer et al., 2017) and Soft-MoE (Puigcerver et al., 2023).
4.4.1. Different numbers of Experts.
We conduct multiple experiments by varying the number of experts in traditional MoE and Soft-MoE, recording the Binary-F1, AUC, and Accuracy for each configuration. The results are summarized in Figure 4.
From the results shown in Figure 4, it is evident that the GMoE consistently outperforms both traditional MoE and Soft-MoE regardless of the number of experts used. Specifically, Soft-MoE performs better than traditional MoE. That is because GMoE uses explicit genre information, eliminating inaccurate dispatch that can occur in traditional MoE and Soft-MoE.
Figure 4. Performance comparison of different numbers of experts in traditional MoE and Soft-MoE. Note that 21 is the number of genres. The results indicate that GMoE outperforms other variants irrespective of the number of experts.
Figure 5. T-SNE visualization of the features processed by GMoE. Different colors represent different genres, indicating distinct clustering of features according to genres.
Table 3. Evaluation of user bias effectiveness. The user bias pre-trained on Dynamic Graph achieves significantly higher accuracy in predicting user labels based on spoiler review percentage compared to raw user embeddings.
feature F1 AUC ACC
DyG pretrained 0.6389 0.7078 0.7777
raw 0.5190 0.5532 0.5874
4.4.2. T-SNE Visualization of Features.
To further validate the effectiveness of the GMoE, we perform dimensionality reduction on the features of all reviews processed by the GMoE layer. Specifically, we first reduce the dimensionality of the features to 50 dimensions using PCA, and then further reduce them to 2 dimensions using T-SNE. The result plot, shown in Figure 5, demonstrates that reviews of different genres exhibit distinct features after being processed by the GMoE.
The T-SNE visualization in Figure 5 indicates that reviews are clustered according to their genres, providing compelling evidence of the effectiveness of our GMoE design. These distinct clusters suggest that GMoE successfully captures and utilizes genre-specific features to enhance its performance. In the visualization, each distinct color represents a different genre. For instance, genres like ’Action’ (red), ’Drama’ (green), and ’Sci-Fi’ (purple) form well-defined clusters, indicating that the features of reviews from these genres are significantly different from each other. The presence of these distinct clusters reaffirms the model’s capability to differentiate and leverage genre-specific information effectively.
Table 4. Examples of the performance of top-2 baselines and GUSD. Underlined parts indicate the plots. ”Key Information” indicates the most helpful information from other sources when detecting spoilers.
Review Text Key Information Label R-GCN MVSD GUSD
(…) The bomb is in a tower in both stories. Glen Manning and Bruce Banner are both there to watch the explosion. Glen and Bruce see some kid in the test area and both dash out to save him, tossing the boy into a ditch instead of jumping in with him. (…) User Bias: Historical Reviews: Review 1: spoiler Review 2: spoiler
Don was an unsuccessful writer. Although he wanted to write a novel, he did not write any word. That should be mainly attributed to his alcoholism. It looked like that without drink he could not live.(…) Movie Genres: Drama Film-Noir
Figure 6. When randomly removing edges from the graph, randomly setting elements of graph features to zero, and reducing the number of labeled samples in the training set, the performance of GUSD gradually declines with these incremental ablations. However, the fact that the model maintains a reasonable level of performance despite these perturbations indicates the robustness of our method.
4.5. User Bias Analysis
To validate whether the obtained user bias truly reflects user preferences, we divide users into two categories based on whether more than 50% of their reviews were spoilers. We then train a simple MLP using user bias as input to predict user labels. By contrast, we also use raw user embeddings that are not trained on a Dynamic Graph as input for predicting user labels. The results, shown in Table 3, demonstrate that the predictions using pretrained user bias are significantly more accurate than those using the raw user embeddings. This indicates that user bias in GUSD effectively captures the users’ preference tendencies.
4.6. Robustness Study
In the realm of social media analysis, particularly for tasks like spoiler detection, obtaining pure and clean data is a significant challenge. Spoiler detection often deals with information that is noisy and heterogeneous. As highlighted by previous studies on bot detection (Cai et al., 2024; Liu et al., 2023), it is crucial to investigate the robustness of models under conditions where information is incomplete or noisy.
To evaluate the robustness of GUSD, we introduce random perturbations to the input data to simulate scenarios where certain information might be missing. Specifically, for the graph structure, we randomly remove some edges; for the graph features, we randomly set some elements to zero; and for the training set, we randomly reduce the number of labeled samples. This experiment aims to assess the model’s performance when parts of the multimodal information are incomplete.
Figure 6 presents the results of our experiments. As shown, even with some missing information, GUSD is still able to make correct predictions in most cases. This supports our hypothesis that multimodal information not only enhances the prediction accuracy of the model but also significantly improves its robustness.
4.7. Case Study
We conduct a case study to explore the impact of movie genres and user bias on spoiler detection. As illustrated in Table 4, our method GUSD, with the guidance of user bias and movie genres, correctly predicts spoilers, whereas the top-2 baseline models (R-GCN and MVSD) fail to do so.
Specifically, in the first case, our GUSD, bolstered by historical review data indicating consistent spoiler-posting behavior from the user, correctly identifies this review as a spoiler. In the second case, based on the Drama and Film-Noir movie genres, GUSD rightly determines that these elements reveal key plot points and hence labels it as a spoiler.
These examples underscore the importance of incorporating user bias and movie genre information. By effectively leveraging such data, GUSD demonstrates superior capability in identifying spoilers compared to baselines.
5. Ethical Considerations
We propose GUSD as an initial filtering tool rather than a final arbiter. Despite attaining state-of-the-art performance, GUSD remains imperfect and must be employed judiciously, in conjunction with human moderators to oversee or halt questionable movie reviews. Additionally, GUSD may reflect the biases embedded in its underlying datasets and models. For example, pretrained language models might perpetuate unwanted social biases and stereotypes (Li et al., 2022; Nadeem et al., 2020). It is within the scope of future research to integrate bias detection and mitigation strategies from machine learning into spoiler detection systems. Furthermore, due to the nature of the task, the dataset includes potentially offensive language, which warrants careful consideration.
6. Conclusion
In this paper, we introduce GUSD, a novel spoiler detection framework that integrates Genreformer and GMoE to effectively model diverse genre features. Additionally, GUSD incorporates dynamic graph pretraining to capture user bias related to spoiler posting. Extensive experiments reveal that GUSD significantly outperforms the state-of-the-art models on two major spoiler detection benchmarks. Further analysis validates the effectiveness of our proposed techniques, demonstrating GUSD’s superior capability in capturing intricate genre features and modeling user bias for spoiler detection.
Appendix A Implementation Details
We use PyTorch (Paszke et al., 2019), scikit-learn (Pedregosa et al., 2011), PyTorch Geometric (Fey and Lenssen, 2019), MoE (Rau, 2019), soft-mixture-of-experts (Puigcerver et al., 2023) and Transformers  (Wolf et al., 2020) library to implement our proposed GUSD. We conduct our experiments on a cluster with 4 Tesla V100 GPUs with 32 GB memory, 16 CPU cores, and 377GB CPU memory. Code, data, hyperparameters and trained models will be made publicly available.
Appendix B Hyperparameters
We present our hyperparameter settings in Table 5 to facilitate reproduction. Code, data, and trained models will be made publicly available.
Appendix C k-hop Neighbors Algorithm
Previous studies (Atwood and Towsley, 2016; Wang et al., 2020) have used the $k$-th power of the adjacency matrix to capture $k$-hop visibility. This method does not strictly define $k$-hop neighbors as nodes reachable in exactly k steps. To address this, we introduce a new GPU-friendly $k$-hop Algorithm that quickly and accurately computes $k$-hop neighbors using adjacency matrices. The details are illustrated in Algorithm 1.
Considering that the number of $k$-hop neighbors $\mathcal{A}_{k}$ ($k>1$) can be large and may require substantial computational resources, we introduce an $r$-ratio (a hyper-parameter) to randomly sample neighbors from the entire set of neighbors for each hop.
Algorithm 1 $k$-hop Neighbors with Minimum Distance
1:Input: adjacency matrix $A$, target $k$-hop number K
2:Output: $k$-hop adjacency matrix $\mathcal{A}_{k}$
3:Initialize distance matrix $D$ with $\infty$, set diagonal to 00
4:Initialize $\mathcal{A}_{1}$ with $A$
5:for all $(i,j)$ do
6:     if $A_{ij}=1$ then
7:         $D_{ij}\leftarrow 1$
8:for $k\leftarrow 2$ to K do
9:     $A^{k}\leftarrow A^{k-1}\times A$
10:     for all $(i,j)$ do
11:         if $A^{k}_{ij}=1$ and $D_{ij}>k$ then
12:              $D_{ij}\leftarrow k$
13:              $(\mathcal{A}_{k})_{ij}\leftarrow 1$
14:return $\mathcal{A}_{k}$
Table 5. Hyperparameter settings of GUSD in Kaggle
Hyperparameter Kaggle LCS
optimizer AdamW AdamW
learning rate 1e-4 1e-4
weight decay 1e-5 1e-5
CrossEntropyLoss weight 1.3 1.3
epochs 20 20
dropout 0.3 0.3
batch size 128 128
early stop patience 3 3
LM Bge-Large (Xiao et al., 2023) Roberta (Liu, 2019)
GNN layer 2 3
GNN input dim 1024 768
GNN hidden dim 1024 768
GNN output dim 1024 768
RetGAT hop num 2 2
RetGAT hop decay $\alpha$ 0.3 0.3
meta dim 3 6
meta encoder hidden dim 1024 768
meta encoder output dim 1024 768
TRM heads 4 4
TRM layers 2 2
GMoE layer 2 2
GMoE hidden dim 1024 768
GMoE output dim 1024 768
GMoE experts number 21 28

### v3

@@
-Unveiling the Hidden: Movie Genre and User Bias in Spoiler Detection
-Abstract.
+
+Abstract
 Spoilers in movie reviews are important on platforms like IMDb and Rotten Tomatoes, offering benefits and drawbacks.
@@
 To address this, we analyze movie review data, finding genre-specific variations in spoiler rates and identifying that certain users are more likely to post spoilers.
-Based on these findings, we introduce a new spoiler detection framework called GUSD  [^1] The code is available at https://github.com/AI-explorer-123/GUSD (Genre-aware and User-specific Spoiler Detection), which incorporates genre-specific data and user behavior bias.
+Based on these findings, we introduce a new spoiler detection framework called GUSD (Genre-aware and User-specific Spoiler Detection), which incorporates genre-specific data and user behavior bias.
 User bias is calculated through dynamic graph modeling of review history.
@@
 This approach advances spoiler detection by addressing genre and user-specific patterns, enhancing user experience on movie review platforms.
-Spoiler Detection, Movie Genre, User Bias, Mixture-of-Experts
-copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
-1. Introduction
-Spoilers in movie reviews have become an important component of the movie-viewing experience on popular platforms like IMDb and Rotten Tomatoes (Cao et al., 2019).
-For those who hope to learn the plot of the movie in advance to judge whether they like it or not, the spoilers are helping, while for those who prefer to experience a movie without prior knowledge, the spoilers can severely diminish the enjoyment by revealing crucial plot points, undermining suspense, and eliciting negative emotions among viewers (Loewenstein, 1994).
+Our source code is available at https://github.com/AI-explorer-123/GUSD
+Keywords:
+Spoiler Detection Movie Genre User Bias Mixture-of-Experts
+1 Introduction
+Spoilers in movie reviews have become an important component of the movie-viewing experience on popular platforms like IMDb and Rotten Tomatoes [5].
+For those who hope to learn the plot of the movie in advance to judge whether they like it or not, the spoilers are helping, while for those who prefer to experience a movie without prior knowledge, the spoilers can severely diminish the enjoyment by revealing crucial plot points, undermining suspense, and eliciting negative emotions among viewers [21].
 Thus, effective spoiler detection methods are crucial for maintaining a positive user experience.
-Figure 1.
-An illustrative example of the data used in our spoiler detection study.
+Figure 1: An illustrative example of the data used in our spoiler detection study.
 The image shows a review of the movie Moana, including the movie’s genres (Adventure, Animation), summary, the review’s content, and user-specific details.
@@
 Existing spoiler detection methods primarily focus on the textual content of reviews.
-For example, DNSD (Chang et al., 2018) integrates review sentences and movie genres, while SpoilerNet (Wan et al., 2019) utilizes a Hierarchical Attention Network and incorporates the item-specificity information.
-More recent approaches, such as MVSD (Wang et al., 2023), incorporate advanced techniques like syntax-aware graph neural networks and external movie knowledge to improve detection performance.
+For example, DNSD [6] integrates review sentences and movie genres, while SpoilerNet [36] utilizes a Hierarchical Attention Network and incorporates the item-specificity information.
+More recent approaches, such as MVSD [38], incorporate advanced techniques like syntax-aware graph neural networks and external movie knowledge to improve detection performance.
 Nevertheless, these methods still exhibit notable limitations.
-Solely relying on textual content proves insufficient for robust spoiler detection (Wang et al., 2023).
+Solely relying on textual content proves insufficient for robust spoiler detection [38].
 Moreover, spoilers are often genre-specific, with varying characteristics depending on the movie’s genre — for instance, suspense films focus on plot details, whereas action films emphasize fight scenes.
@@
 Leveraging these user-specific behavior bias can improve the detection performance by allowing models to adapt to user behavior bias.
-Figure 2. (a) Spoiler rate across different movie genres (partial) in LCS dataset. (b) Kernel density estimation plot and distribution histogram of spoiler across different users.
+Figure 2: (a) Spoiler rate across different movie genres (partial) in LCS dataset. (b) Kernel density estimation plot and distribution histogram of spoiler across different users.
 To address the challenges of genre-specific spoiler tendencies and user bias in spoiler detection, we propose a comprehensive framework named GUSD (Genre-aware and User-specific Spoiler Detection).
@@
 Extensive experiments on two benchmark datasets demonstrate its robustness and effectiveness, showing superior performance across various conditions.
-Figure 3.
-Overview of our proposed GUSD framework, which integrates genre-specific information, user behavior bias, and global perceptive RetGAT for spoiler detection.
+2 Related Work
+2.1 Spoiler Detection
+The goal of automatic spoiler detection is to identify spoilers in reviews from domains like television [3], books [36], and movies [3].
+Existing approaches to spoiler detection can be broadly classified into three categories: keyword matching methods, machine learning techniques, and deep learning models.
+2.1.1 Keyword matching methods.
+These approaches rely on a set of predefined keywords to identify spoilers.
+Examples include keywords related to sports teams or events [25], or actors’ names [12].
+Although useful in specific scenarios, this method requires manual keyword definition and lacks generalizability across different application contexts.
+2.1.2 Machine learning techniques.
+These methods often involve topic modeling or support vector machines using handcrafted features.
+For example, Guo et al.
+[13] applied a bag-of-words representation combined with an LDA-based model for spoiler detection.
+Jeon et al.
+[16] developed an SVM classifier incorporating four extracted features, while Boyd et al.
+[3] utilized lexical features and meta-data of review subjects (e.g., movies and books) in an SVM model.
+2.1.3 Deep learning models.
+These models mainly leverage NLP techniques, employing RNNs, LSTMs, Transformer, and language models to process review texts and movie information through end-to-end training.
+Bao et al.
+[2] utilized LSTMs, BERT, and RoBERTa for sentence-level spoiler detection.
+DNSD [6] focused on incorporating external genre information using GRU and CNN.
+SpoilerNet [36] introduced item-specificity and bias with bi-RNN enhanced by GRU.
+SDGNN [7] leveraged dependency relations between context words in sentences with graph neural networks to capture semantics.
+While some existing methods incorporate genre information and user bias  [6, 40], they often rely on quite simple techniques such as concatenating or adding these additional features to the initial review features.
+Such approaches lack the sophistication needed for more effective and intricate modeling of genre features and user biases.
+2.2 Mixture of Experts
+The Mixture of Experts (MoE) approach, grounded in the Divide-and-Conquer principle, segments an input sample into sub-tasks and trains specialized experts for each sub-task.
+This method is extensively utilized in NLP to boost model capacity [33] and enhance reasoning capabilities [23].
+Shazeer et al.
+[33] introduced a sparsely-gated Mixture-of-Experts layer, enabling conditional computing in large language models.
+Fedus et al.
+[10] developed simplified routing algorithms for MoE to enhance training stability and reduce computational costs.
+Furthermore, Soft-MoE [29] was introduced to mitigate issues like training instability and token dropping inherent in traditional MoE approaches.
+Despite these advancements, traditional MoE methods assign tokens dynamically, which can cause incorrect token assignment, particularly when the dataset contains explicit category information such as movie genres and their associated reviews.
+Figure 3: Overview of our proposed GUSD framework, which integrates genre-specific information, user behavior bias, and global perceptive RetGAT for spoiler detection.
 It preprocesses movie, user, and review data with MLP and language models, and captures user bias via dynamic graph modeling.
@@
 Finally, a classifier detects spoilers using the aggregated features.
-2. Related Work
-2.1. Spoiler Detection
-The goal of automatic spoiler detection is to identify spoilers in reviews from domains like television (Boyd-Graber et al., 2013), books (Wan et al., 2019), and movies (Boyd-Graber et al., 2013).
-Existing approaches to spoiler detection can be broadly classified into three categories: keyword matching methods, machine learning techniques, and deep learning models.
-2.1.1. Keyword matching methods.
-These approaches rely on a set of predefined keywords to identify spoilers.
-Examples include keywords related to sports teams or events (Nakamura and Tanaka, 2007), or actors’ names (Golbeck, 2012).
-Although useful in specific scenarios, this method requires manual keyword definition and lacks generalizability across different application contexts.
-2.1.2. Machine learning techniques.
-These methods often involve topic modeling or support vector machines using handcrafted features.
-For example, Guo et al.
-(Guo and Ramakrishnan, 2010) applied a bag-of-words representation combined with an LDA-based model for spoiler detection.
-Jeon et al.
-(Jeon et al., 2013) developed an SVM classifier incorporating four extracted features, while Boyd et al.
-(Boyd-Graber et al., 2013) utilized lexical features and meta-data of review subjects (e.g., movies and books) in an SVM model.
-2.1.3. Deep learning models.
-These models mainly leverage NLP techniques, employing RNNs, LSTMs, Transformer, and language models to process review texts and movie information through end-to-end training.
-Bao et al.
-(Bao et al., 2021) utilized LSTMs, BERT, and RoBERTa for sentence-level spoiler detection.
-DNSD (Chang et al., 2018) focused on incorporating external genre information using GRU and CNN.
-SpoilerNet (Wan et al., 2019) introduced item-specificity and bias with bi-RNN enhanced by GRU.
-SDGNN (Chang et al., 2021) leveraged dependency relations between context words in sentences with graph neural networks to capture semantics.
-While some existing methods incorporate genre information and user bias  (Chang et al., 2018; Wróblewska et al., 2021), they often rely on quite simple techniques such as concatenating or adding these additional features to the initial review features.
-Such approaches lack the sophistication needed for more effective and intricate modeling of genre features and user biases.
-2.2. Mixture of Experts
-The Mixture of Experts (MoE) approach, grounded in the Divide-and-Conquer principle, segments an input sample into sub-tasks and trains specialized experts for each sub-task.
-This method is extensively utilized in NLP to boost model capacity (Shazeer et al., 2017) and enhance reasoning capabilities (Madaan et al., 2021).
-Shazeer et al.
-(Shazeer et al., 2017) introduced a sparsely-gated Mixture-of-Experts layer, enabling conditional computing in large language models.
-Fedus et al.
-(Fedus et al., 2022) developed simplified routing algorithms for MoE to enhance training stability and reduce computational costs.
-Furthermore, Soft-MoE (Puigcerver et al., 2023) was introduced to mitigate issues like training instability and token dropping inherent in traditional MoE approaches.
-Despite these advancements, traditional MoE methods assign tokens dynamically, which can cause incorrect token assignment, particularly when the dataset contains explicit category information such as movie genres and their associated reviews.
-3. Methodology
+3 Methodology
 Figure 3 shows the architecture of our proposed GUSD framework.
@@
 Finally, a classifier utilizes the aggregated features to detect spoilers.
-3.1. Data preprocessing
-3.1.1. Meta information.
-For review nodes, user nodes, and movie nodes, each type possesses meta information.
-After collecting the meta information for all three types of nodes, we pad them to the same length.
+3.1 Data preprocessing
+3.1.1 Meta information.
+For review nodes, user nodes, and movie nodes, each type possesses metadata (details can be found in the supplementary material).
+After collecting the metadata for all three types of nodes, we pad them to the same length.
 A two-layer MLP is then employed as the meta encoder, producing the meta embeddings $N_{m}$.
-3.1.2. Textual information.
+3.1.2 Textual information.
 The textual content is fundamental for effective spoiler detection.
@@
 The encoded embeddings are then transformed using a single-layer MLP, producing refined embeddings $N_{t}$.
-3.1.3. User bias acquisition using dynamic-graph pre-training.
+3.1.3 User bias acquisition using dynamic-graph pre-training.
 To better capture the dynamic information attributes of users, we adopt the dynamic graph to handle users’ review history flexibly.
@@
 The details about the formation of the graph will be displayed in Section 3.2.
-Considering the absence of additional information about the dynamic edges, we simply initial the edge features as zero vectors.
-Then, we employ the robust DyGFormer (Yu et al., 2023) as our dynamic graph encoder to capture the interactions among various nodes and utilize Link Prediction as the downstream task.
+Considering the absence of additional information about the dynamic edges, we simply initialize the edge features as zero vectors.
+Then, we employ the robust DyGFormer [42] as our dynamic graph encoder to capture the interactions among various nodes and utilize Link Prediction as the downstream task.
 We specifically obtain the features of user nodes as user bias $U_{b}$ from the dynamic graph encoder.
-3.2. R2GFormer
+3.2 R2GFormer
 After acquiring the initial meta and textual features of the nodes, we use the textual features $N_{t}$, which include rich information about the nodes, as the initial embeddings $N_{g}^{(0)}$ for R2GFormer.
@@
 Then, we decompose an R2GFormer layer into two parts: RetGAT and Genreformer, which will be introduced respectively.
-3.2.1. Graph Construction.
+3.2.1 Graph Construction.
 We first construct a directed graph consisting of three types of nodes: { User, Review, Movie } and the following three types of edges:
@@
 E3: User-Review We connect a user node to a review node if the user posts the review.
-3.2.2. RetGAT.
-Inspired by the work of (Sun et al., 2023; Fan et al., 2024; Nikolentzos et al., 2020), we propose RetGAT, which extends the RetNet framework by integrating a global perception into GAT, incorporating explicit exponential decay for nodes within $k$ hops and truncation for nodes beyond $k$ hops.
+3.2.2 RetGAT.
+Inspired by the work of [34, 9, 26], we propose RetGAT, which extends the RetNet framework by integrating a global perception into GAT, incorporating explicit exponential decay for nodes within $k$ hops and truncation for nodes beyond $k$ hops.
 This method ensures a broad receptive field while balancing performance and computational complexity by dynamically adjusting the influence of nodes based on their distance, with closer nodes having a higher impact on the final node features.
@@
 Beyond $k$ hops, the influence of nodes is truncated to maintain computational efficiency and focus on relevant information within the $k$-hop range.
-Note that we correct the algorithm that previous work (Atwood and Towsley, 2016; Wang et al., 2020) used to compute $k$-hop neighbors, and the details are available in the supplementary material.
+Note that we correct the algorithm that previous work [1, 37] used to compute $k$-hop neighbors, and the details are available in the supplementary material.
 The decay factor $\delta_{h}$ for hop $h$ ($h\leq k$) is defined as:
-(1) $\delta_{h}$ $=\exp(-\alpha\cdot h),$
+$\delta_{h}$ $=\exp(-\alpha\cdot h),$ (1)
 where $\alpha$ is a hyperparameter controlling the intensity of exponential decay. For the $l$-th layer, the contribution of the $h$-th hop neighbors is computed as:
@@
 The final node features ${N_{g}}^{(l)}$ for layer $l$ are aggregated from all the $k$-hop contributions using an Aggregator function:
-(3) ${N_{g}^{(l)}}$ $=\text{AGGREGATOR}_{r}^{(l)}(\text{${N_{g,1}^{(l)}}$},\text{${N_{g,2}^{(l)}}$},\ldots,\text{${N_{g,h}^{(l)}}$},\ldots,\text{${N_{g,k}^{(l)}}$}).$
+${N_{g}^{(l)}}$ $=\text{AGGREGATOR}_{r}^{(l)}(\text{${N_{g,1}^{(l)}}$},\text{${N_{g,2}^{(l)}}$},\ldots,\text{${N_{g,h}^{(l)}}$},\ldots,\text{${N_{g,k}^{(l)}}$}).$ (3)
 where $\text{AGGREGATOR}_{r}^{(l)}$ is the aggregator for RetGAT at the l-th layer, which can be summation, concatenation, or a TransformerEncoder (TRM).
-3.2.3. Genreformer.
+3.2.3 Genreformer.
 After message passing among different $k$-hop neighbors, Genreformer aims to pay attention to the entire graph to extract more comprehensive features of different genres.
 First, the global genre-specific representation is obtained by aggregating the features of all the review and movie nodes belonging to the same genre in the $l$-th layer, which is as follows:
-(4) $g_{j}^{(l)}=\text{AGGREGATOR}^{(l)}_{g}(\{n_{g,i}^{(l)}\mid i\in\mathcal{N}_{j}\}),$
+$g_{j}^{(l)}=\text{AGGREGATOR}^{(l)}_{g}(\{n_{g,i}^{(l)}\mid i\in\mathcal{N}_{j}\}),$ (4)
 where $n_{g,i}^{(l)}$ is the feature of node $i$ in the $l$-th layer, $\mathcal{N}_{j}$ denotes the set of nodes (both review and movie nodes) belonging to the $j$-th genre, and $\text{AGGREGATOR}^{(l)}_{g}$ is the aggregator of Genreformer in the $l$-th layer, which can be summation, concatenation, or a TRM.
 Next, we use a TRM to facilitate inter-genre information interaction, allowing each genre to acquire information from similar genres to further enrich its features:
-(5) $[g_{1}^{(l)}g_{2}^{(l)}\cdots g_{j}^{(l)}\cdots g_{c}^{(l)}]=\text{TRM}([g_{1}^{(l)}g_{2}^{(l)}\cdots g_{j}^{(l)}\cdots g_{c}^{(l)}]),$
+$[g_{1}^{(l)}g_{2}^{(l)}\cdots g_{j}^{(l)}\cdots g_{c}^{(l)}]=\text{TRM}([g_{1}^{(l)}g_{2}^{(l)}\cdots g_{j}^{(l)}\cdots g_{c}^{(l)}]),$ (5)
 where $c$ is the number of genres.
@@
 Subsequently, we concatenate the review or movie feature and its genre feature, then use an MLP to project it into the desired feature space, $i.e.$,
-(6) $z_{i}^{(l)}$ $=\text{MEAN}(\{g_{j}^{(l)}\mid j\in G_{i}\}),$
-(7) $n_{g,i}^{(l)}$ $=\text{\text{MLP}}([n_{g,i}^{(l)}\parallel z_{i}^{(l)}]).$
+$z_{i}^{(l)}$ $=\text{MEAN}(\{g_{j}^{(l)}\mid j\in G_{i}\}),$ (6)
+$n_{g,i}^{(l)}$ $=\text{\text{MLP}}([n_{g,i}^{(l)}\parallel z_{i}^{(l)}]).$ (7)
 where $G_{i}$ denotes the set of genres to which node $i$ belongs, $n_{g,i}^{(l)}$ denotes the feature of node $i$ in the $l$-th layer, $z_{i}^{(l)}$ is the aggregated genre feature for node $i$ in the $l$-th layer, and $\parallel$ means concatenation.
 The MLP projects the concatenated feature into the desired feature space.
-3.2.4. Overall interaction.
+3.2.4 Overall interaction.
 One layer of our proposed R2GFormer layer, however, cannot enable the information interaction between all information sources.
@@
 This process can be formulated as follows:
-(8) $N_{g}^{(l)}=\text{R2GFormer}(N_{g}^{(l-1)},\mathcal{E},G).$
+$N_{g}^{(l)}=\text{R2GFormer}(N_{g}^{(l-1)},\mathcal{E},G).$ (8)
 where $\mathcal{E}$ denotes all the edges of the sampling graph, $G$ denotes the genres of every movie and review node.
 After $L\times$ R2GFormer layers, we obtain the final graph representation $N_{g}^{(L)}$ of all nodes.
-3.3. Multimodal Fusion
+3.3 Multimodal Fusion
 Through the data above processed by the R2GFormer, we have obtained the graph structural information and meta information for all nodes, including user, review and movie nodes.
@@
 For each type of node, we utilize a type-wise TRM to facilitate inter-modal information interaction, then concatenate features of different modals followed by an MLP to get the final representation for each node, i.e.,
-(9) $U_{g},U_{m},U_{b}=\text{TRM}([U_{g}\ U_{m}\ U_{b}]$ $,U=\text{\text{MLP}}([U_{g}\parallel U_{m}\parallel U_{b}])),$
-(10) $R_{g},R_{m}=\text{TRM}([R_{g}\ R_{m}]$ $,R=\text{\text{MLP}}([R_{g}\parallel R_{m}])),$
-(11) $M_{g},M_{m}=\text{TRM}([M_{g}\ M_{m}]$ $,M=\text{\text{MLP}}([M_{g}\parallel M_{m}]),$
+$U_{g},U_{m},U_{b}=\text{TRM}([U_{g}\ U_{m}\ U_{b}]$ $,U=\text{\text{MLP}}([U_{g}\parallel U_{m}\parallel U_{b}])),$ (9)
+$R_{g},R_{m}=\text{TRM}([R_{g}\ R_{m}]$ $,R=\text{\text{MLP}}([R_{g}\parallel R_{m}])),$ (10)
+$M_{g},M_{m}=\text{TRM}([M_{g}\ M_{m}]$ $,M=\text{\text{MLP}}([M_{g}\parallel M_{m}]),$ (11)
 where $U_{g}$, $R_{g}$ and $M_{g}$ are derived from $N_{g}$, $U_{m}$, $R_{m}$ and $M_{m}$ are dervied from $N_{m}$.
 After obtaining the comprehensive representation of each type of node, we then concatenate each review feature $r_{i}$ with its corresponding movie feature $m_{i}$ and user feature $u_{i}$:
-(12) $r_{i}=[\ r_{i}\parallel m_{i}\parallel u_{i}\ ].$
-3.4. GMoE
+$r_{i}=[\ r_{i}\parallel m_{i}\parallel u_{i}\ ].$ (12)
+3.4 GMoE
 Inspired by the successful applications of Mixture-of-Experts in NLP and bot detection, and its capability to handle the small subsets of the whole dataset, we adopt MoE to handle different genres of reviews.
@@
 Specifically, instead of using the gating mechanism in the traditional MoE structure, we assign tokens to experts simply according to their genres: which genre it belongs to, which expert will deal with it; how many genres it belongs to, how many experts will deal with it.
-(13) $r_{i}=\text{AGGREGATOR}_{m}(\{\text{Expert}_{j}(r_{i})\mid\forall{j}\in{G_{i}}\}).$
+$r_{i}=\text{AGGREGATOR}_{m}(\{\text{Expert}_{j}(r_{i})\mid\forall{j}\in{G_{i}}\}).$ (13)
 where $G_{i}$ denotes the set of genres to which node $i$ belongs; AGGRE-
 GATORm can be summation, concatenation, or a TRM; each Expert is a MLP for simplicity.
-3.5. Learning and Optimization
+3.5 Learning and Optimization
 After using GMoE to process genre-specific information, we acquire the final representation $r_{i}$ for the $i$-th review.
@@
 The total loss function is as follows:
-(14) $Loss=-\sum_{i\in\mathcal{R}}{y_{i}\log\hat{y_{i}}}+\lambda\sum_{\theta\in\Theta}{\theta^{2}}.$
+$Loss=-\sum_{i\in\mathcal{R}}{y_{i}\log\hat{y_{i}}}+\lambda\sum_{\theta\in\Theta}{\theta^{2}}.$ (14)
 where $\hat{y_{i}}$ and $y_{i}$ are the prediction for the $i$-th review and its corresponding ground truth, respectively.
 $\mathcal{R}$ encompasses all the reviews in the training set, while $\Theta$ denotes all trainable model parameters in GUSD, and $\lambda$ is a hyper-parameter that maintains the balance between the two parts.
-Table 1.
-Accuracy, AUC, and binary F1-score of GUSD and three types of baseline methods on two spoiler detection datasets.
+4 Experiment
+4.1 Experiment Settings
+4.1.1 Dataset.
+To evaluate our GUSD framework along with 14 other representative baselines on two widely recognized datasets: LCS [38] and Kaggle [24]:
+  • LCS is a comprehensive dataset for automatic spoiler detection, comprising 1,860,715 reviews, 259,705 users, and 147,191 movies. And about 24.59% (457,500) of the reviews are spoilers.
+  • Kaggle, introduced in 2019, consists of 573,913 valid reviews, 263,407 users, and 1,572 movies. And about 25.87% (150,924) of the reviews are spoilers.
+Note that both datasets include the genre information of all movies, with specific categories defined according to IMDb standards, which facilitates the operation of our GUSD framework.
+Following MVSD [38], we randomly split the reviews into training, validation, and test sets with a ratio of 7:2:1.
+4.1.2 Baselines.
+To achieve a comprehensive evaluation, we compare GUSD with pre-trained language models, GNN-based models, and task-specific baselines.
+For the pre-trained language models, the procedure involves feeding the review text into the model, averaging all token embeddings, and then applying two fully connected layers to perform spoiler detection.
+Regarding the GNN-based models, the graph neural network takes the output of RoBERTa [19] as the initial node features.
+Below, we provide a concise overview of each baseline method.
+  • BERT [8] is a language model pre-trained on extensive natural language data, using masked language modeling and next sentence prediction tasks.
+  • RoBERTa [19] improves upon BERT by eliminating the next sentence prediction task and enhancing masking techniques.
+  • BART [18] is a pre-trained language model that advances traditional autoregressive models through bidirectional encoding and denoising objectives.
+  • DeBERTa [14] refines BERT by implementing disentangled attention and an improved mask decoder, making it a more advanced language model.
+  • Bge-Large [41] is trained on a comprehensive training dataset C-MTP, combining vast unlabeled data and diverse labeled data.
+  • GCN [17] is a foundational graph neural network that performs convolutions on graph nodes and their neighbors, effectively propagating information.
+  • R-GCN [32] extends GCN to handle multi-relational graphs by incorporating relation-specific weights.
+  • GAT [35] is a graph neural network that applies attention mechanisms to dynamically assign importance to neighboring nodes.
+  • SimpleHGN [22] is tailored for heterogeneous graphs, integrating multiple types of nodes and edges with a shared embedding space and adaptive aggregation strategies.
+  • GPS [30] propose a recipe to build a general, powerful, scalable graph Transformer with linear complexity.
+  • HGT [15] design node- and edge-type dependent parameters to characterize the heterogeneous attention over each edge for modeling Web-scale heterogeneous graphs.
+  • DNSD [6] is a spoiler detection method that employs a CNN-based genre-aware attention mechanism.
+  • SpoilerNet [36] uses a hierarchical attention network and GRU alongside item and user bias terms for spoiler detection.
+  • MVSD [38] leverages external movie knowledge and user networks to detect spoilers.
+Table 1: Accuracy, AUC, and binary F1-score of GUSD and three types of baseline methods on two spoiler detection datasets.
 We run all experiments five times to ensure a consistent evaluation and report the average performance as well as standard deviation in parentheses.
@@
 Ours 80.24 $\pm$0.73 87.00 $\pm$0.37 89.65 $\pm$0.36 75.37 $\pm$0.10 83.71 $\pm$0.27 88.32 $\pm$0.08
-4. Experiment
-4.1. Experiment Settings
-4.1.1. Dataset.
-To evaluate our GUSD framework along with 14 other representative baselines on two widely recognized datasets: LCS (Wang et al., 2023) and Kaggle (Wan et al., 2019):
-  • LCS is a comprehensive dataset for automatic spoiler detection, comprising 1,860,715 reviews, 259,705 users, and 147,191 movies. And about 24.59% (457,500) of the reviews are spoilers.
-  • Kaggle, introduced in 2019, consists of 573,913 valid reviews, 293,407 users, and 1,572 movies. And about 25.87% (150,924) of the reviews are spoilers.
-Note that both datasets include the genre information of all movies, with specific categories defined according to IMDb standards, which facilitates the operation of our GUSD framework.
-Following MVSD (Wang et al., 2023), we randomly split the reviews into training, validation, and test sets with a ratio of 7:2:1.
-4.1.2. Baselines.
-To achieve a comprehensive evaluation, we compare GUSD with pre-trained language models, GNN-based models, and task-specific baselines.
-For the pre-trained language models, the procedure involves feeding the review text into the model, averaging all token embeddings, and then applying two fully connected layers to perform spoiler detection.
-Regarding the GNN-based models, the graph neural network takes the output of RoBERTa (Liu, 2019) as the initial node features.
-Below, we provide a concise overview of each baseline method.
-  • BERT (Devlin, 2018) is a language model pre-trained on extensive natural language data, using masked language modeling and next sentence prediction tasks.
-  • RoBERTa (Liu, 2019) improves upon BERT by eliminating the next sentence prediction task and enhancing masking techniques.
-  • BART (Lewis, 2019) is a pre-trained language model that advances traditional autoregressive models through bidirectional encoding and denoising objectives.
-  • DeBERTa (He et al., 2021) refines BERT by implementing disentangled attention and an improved mask decoder, making it a more advanced language model.
-  • Bge-Large (Xiao et al., 2023) is trained on a comprehensive training dataset C-MTP, combining vast unlabeled data and diverse labeled data.
-  • GCN (Kipf and Welling, 2016) is a foundational graph neural network that performs convolutions on graph nodes and their neighbors, effectively propagating information.
-  • R-GCN (Schlichtkrull et al., 2018) extends GCN to handle multi-relational graphs by incorporating relation-specific weights.
-  • GAT (Veličković et al., 2017) is a graph neural network that applies attention mechanisms to dynamically assign importance to neighboring nodes.
-  • SimpleHGN (Lv et al., 2021) is tailored for heterogeneous graphs, integrating multiple types of nodes and edges with a shared embedding space and adaptive aggregation strategies.
-  • GPS (Rampášek et al., 2022) propose a recipe to build a general, powerful, scalable graph Transformer with linear complexity.
-  • HGT (Hu et al., 2020) design node- and edge-type dependent parameters to characterize the heterogeneous attention over each edge for modeling Web-scale heterogeneous graphs.
-  • DNSD (Chang et al., 2018) is a spoiler detection method that employs a CNN-based genre-aware attention mechanism.
-  • SpoilerNet (Wan et al., 2019) uses a hierarchical attention network and GRU alongside item and user bias terms for spoiler detection.
-  • MVSD (Wang et al., 2023) leverages external movie knowledge and user networks to detect spoilers.
-4.2. Main Results
+4.2 Main Results
 We evaluated our GUSD framework and 14 other baselines on two datasets. The results presented in Table 1 demonstrate the following:
   • GUSD consistently outperforms all baselines across both datasets.
-Specifically, compared with the previous state-of-the-art method MVSD (Wang et al., 2023), GUSD achieves 6.1% higher Binary-F1, 5.5% higher AUC, and 2.0% higher accuracy on the LCS dataset, as well as 15.2% higher Binary-F1, 11.6% higher AUC, and 6.1% higher accuracy on the Kaggle dataset.
+Specifically, compared with the previous state-of-the-art method MVSD [38], GUSD achieves 6.1% higher Binary-F1, 5.5% higher AUC, and 2.0% higher accuracy on the LCS dataset, as well as 15.2% higher Binary-F1, 11.6% higher AUC, and 6.1% higher accuracy on the Kaggle dataset.
 These improvements are statistically significant.
   • In both datasets, graph-based models generally outperform other types of baselines, reaffirming the importance of analyzing the graph structure of reviews and their corresponding users and movies.
-  • Compared with DNSD (Chang et al., 2018), which also focuses on genre features of the reviews, GUSD surpasses DNSD across all three metrics in both datasets, further proving the effectiveness and robustness of our global-aware genreformer and GMoE methods.
-  • Both SpoilerNet (Wan et al., 2019) and GUSD utilize user bias, but GUSD outperforms SpoilerNet in all three metrics across both datasets.
+  • Compared with DNSD [6], which also focuses on genre features of the reviews, GUSD surpasses DNSD across all three metrics in both datasets, further proving the effectiveness and robustness of our global-aware genreformer and GMoE methods.
+  • Both SpoilerNet [36] and GUSD utilize user bias, but GUSD outperforms SpoilerNet in all three metrics across both datasets.
 This demonstrates that our dynamic graph pretraining can better identify the latent behavior pattern of whether a user is likely to post spoilers.
-4.3. Ablation Study
+4.3 Ablation Study
 As GUSD outperforms all the baselines and has reached state-of-the-art (SOTA) performance across the two datasets, we conducted ablation study to further explore the impact of each part of GUSD on the final performance with the Kaggle Dataset.
@@
 The results in Table 2 show an obvious decrease in performance, confirming that user bias information is critical for effective spoiler detection.
-  • For the RetGAT component, we evaluated two variations: using an approximate method to compute $k$-hop neighbors (Atwood and Towsley, 2016; Wang et al., 2020) and replacing our RetGAT with a standard GAT.
+  • For the RetGAT component, we evaluated two variations: using an approximate method to compute $k$-hop neighbors [1, 37] and replacing our RetGAT with a standard GAT.
 Both variations lead to a drop in performance, indicating the necessity of our RetGAT design for capturing appropriate graph structures.
-  • To investigate the impact of GMoE, we replace it with a simple MLP, a traditional MoE (Shazeer et al., 2017), and a Soft-MoE (Puigcerver et al., 2023).
+  • To investigate the impact of GMoE, we replace it with a simple MLP, a traditional MoE [33], and a Soft-MoE [29].
 Note that we set the number of experts of the traditional MoE and Soft-MoE to be the same as in GMoE, i.e., the number of genres.
@@
 The results in Table 2 show that the full Genreformer with sum AGGREGATOR outperforms these ablated versions, validating the necessity of the Genreformer.
-Table 2. Ablation study of GUSD on Kaggle Dataset. Bold indicates the best performance, underline the second best.
+Table 2: Ablation study of GUSD on Kaggle Dataset. Bold indicates the best performance, underline the second best.
 Catagory Ablation Settings F1 AUC Acc
@@
 GMoE MLP 78.52 86.09 88.25
-MoE (Shazeer et al., 2017) 78.75 86.36 88.54
-Soft MoE (Puigcerver et al., 2023) 79.10 86.10 88.78
+MoE [33] 78.75 86.36 88.54
+Soft MoE [29] 79.10 86.10 88.78
 Genreformer -w/o genreformer 77.73 84.68 88.13
@@
 Ours GUSD 80.24 87.00 89.65
-4.4. GMoE Study
-In this section, we conduct further experiments on the GMoE to better understand its effectiveness compared to other previous MoE methods such as traditional MoE (Shazeer et al., 2017) and Soft-MoE (Puigcerver et al., 2023).
-4.4.1. Different numbers of Experts.
+4.4 GMoE Study
+In this section, we conduct further experiments on the GMoE to better understand its effectiveness compared to other previous MoE methods such as traditional MoE [33] and Soft-MoE [29].
+4.4.1 Different numbers of Experts.
 We conduct multiple experiments by varying the number of experts in traditional MoE and Soft-MoE, recording the Binary-F1, AUC, and Accuracy for each configuration.
@@
 That is because GMoE uses explicit genre information, eliminating inaccurate dispatch that can occur in traditional MoE and Soft-MoE.
-Figure 4.
-Performance comparison of different numbers of experts in traditional MoE and Soft-MoE.
+Figure 4: Performance comparison of different numbers of experts in traditional MoE and Soft-MoE.
 Note that 21 is the number of genres.
 The results indicate that GMoE outperforms other variants irrespective of the number of experts.
-Figure 5. T-SNE visualization of the features processed by GMoE. Different colors represent different genres, indicating distinct clustering of features according to genres.
-Table 3.
-Evaluation of user bias effectiveness.
-The user bias pre-trained on Dynamic Graph achieves significantly higher accuracy in predicting user labels based on spoiler review percentage compared to raw user embeddings.
-feature F1 AUC ACC
-DyG pretrained 0.6389 0.7078 0.7777
-raw 0.5190 0.5532 0.5874
-4.4.2. T-SNE Visualization of Features.
+Figure 5: T-SNE visualization of the features processed by GMoE. Different colors represent different genres, indicating distinct clustering of features according to genres.
+4.4.2 T-SNE Visualization of Features.
 To further validate the effectiveness of the GMoE, we perform dimensionality reduction on the features of all reviews processed by the GMoE layer.
@@
 The presence of these distinct clusters reaffirms the model’s capability to differentiate and leverage genre-specific information effectively.
-Table 4.
-Examples of the performance of top-2 baselines and GUSD.
+5 Conclusion
+In this paper, we introduce GUSD, a novel spoiler detection framework that integrates Genreformer and GMoE to effectively model diverse genre features.
+Additionally, GUSD incorporates dynamic graph pretraining to capture user bias related to spoiler posting.
+Extensive experiments reveal that GUSD significantly outperforms the state-of-the-art models on two major spoiler detection benchmarks.
+Further analysis validates the effectiveness of our proposed techniques, demonstrating GUSD’s superior capability in capturing intricate genre features and modeling user bias for spoiler detection.
+{credits}
+5.0.1 Acknowledgements
+This work was supported by the National Nature Science Foundation of China (No.
+62272374, No.
+62192781 ), the Natural Science Foundation of Shaanxi Province (2024JC-JCQN-62), the National Nature Science Foundation of China (No.
+62202367, No.
+62250009), the Key Research and Development Project in Shaanxi Province No.
+2023GXLH-024, Project of China Knowledge Center for Engineering Science and Technology, and Project of Chinese academy of engineering “The Online and Offline Mixed Educational Service System for ‘The Belt and Road’ Training in MOOC China”, and the K.
+C.
+Wong Education Foundation.
+Lastly, we would like to thank all LUD lab members for fostering a collaborative research environment.
+Appendix 0.A Implementation Details
+We use PyTorch [27], scikit-learn [28], PyTorch Geometric [11], MoE [31], soft-mixture-of-experts [29] and Transformers  [39] library to implement our proposed GUSD.
+We conduct our experiments on a cluster with 4 Tesla V100 GPUs with 32 GB memory, 16 CPU cores, and 377GB CPU memory.
+Code, data, hyperparameters and trained models will be made publicly available.
+Appendix 0.B More Experiments
+Table 3: Evaluation of user bias effectiveness.
+The user bias pre-trained on Dynamic Graph achieves significantly higher accuracy in predicting user labels based on spoiler review percentage compared to raw user embeddings.
+feature F1 AUC ACC
+DyG pretrained 0.6389 0.7078 0.7777
+raw 0.5190 0.5532 0.5874
+Table 4: Examples of the performance of top-2 baselines and GUSD.
 Underlined parts indicate the plots.
-”Key Information” indicates the most helpful information from other sources when detecting spoilers.
+"Key Information" indicates the most helpful information from other sources when detecting spoilers.
 Review Text Key Information Label R-GCN MVSD GUSD
@@
 It looked like that without drink he could not live.(…) Movie Genres: Drama Film-Noir
-Figure 6.
-When randomly removing edges from the graph, randomly setting elements of graph features to zero, and reducing the number of labeled samples in the training set, the performance of GUSD gradually declines with these incremental ablations.
+Figure 6: When randomly removing edges from the graph, randomly setting elements of graph features to zero, and reducing the number of labeled samples in the training set, the performance of GUSD gradually declines with these incremental ablations.
 However, the fact that the model maintains a reasonable level of performance despite these perturbations indicates the robustness of our method.
-4.5. User Bias Analysis
+0.B.1 User Bias Analysis
 To validate whether the obtained user bias truly reflects user preferences, we divide users into two categories based on whether more than 50% of their reviews were spoilers.
@@
 This indicates that user bias in GUSD effectively captures the users’ preference tendencies.
-4.6. Robustness Study
+0.B.2 Robustness Study
 In the realm of social media analysis, particularly for tasks like spoiler detection, obtaining pure and clean data is a significant challenge.
 Spoiler detection often deals with information that is noisy and heterogeneous.
-As highlighted by previous studies on bot detection (Cai et al., 2024; Liu et al., 2023), it is crucial to investigate the robustness of models under conditions where information is incomplete or noisy.
+As highlighted by previous studies on bot detection [4, 20], it is crucial to investigate the robustness of models under conditions where information is incomplete or noisy.
 To evaluate the robustness of GUSD, we introduce random perturbations to the input data to simulate scenarios where certain information might be missing.
@@
 This supports our hypothesis that multimodal information not only enhances the prediction accuracy of the model but also significantly improves its robustness.
-4.7. Case Study
+0.B.3 Case Study
 We conduct a case study to explore the impact of movie genres and user bias on spoiler detection.
@@
 By effectively leveraging such data, GUSD demonstrates superior capability in identifying spoilers compared to baselines.
-5. Ethical Considerations
-We propose GUSD as an initial filtering tool rather than a final arbiter.
-Despite attaining state-of-the-art performance, GUSD remains imperfect and must be employed judiciously, in conjunction with human moderators to oversee or halt questionable movie reviews.
-Additionally, GUSD may reflect the biases embedded in its underlying datasets and models.
-For example, pretrained language models might perpetuate unwanted social biases and stereotypes (Li et al., 2022; Nadeem et al., 2020).
-It is within the scope of future research to integrate bias detection and mitigation strategies from machine learning into spoiler detection systems.
-Furthermore, due to the nature of the task, the dataset includes potentially offensive language, which warrants careful consideration.
-6. Conclusion
-In this paper, we introduce GUSD, a novel spoiler detection framework that integrates Genreformer and GMoE to effectively model diverse genre features.
-Additionally, GUSD incorporates dynamic graph pretraining to capture user bias related to spoiler posting.
-Extensive experiments reveal that GUSD significantly outperforms the state-of-the-art models on two major spoiler detection benchmarks.
-Further analysis validates the effectiveness of our proposed techniques, demonstrating GUSD’s superior capability in capturing intricate genre features and modeling user bias for spoiler detection.
-Appendix A Implementation Details
-We use PyTorch (Paszke et al., 2019), scikit-learn (Pedregosa et al., 2011), PyTorch Geometric (Fey and Lenssen, 2019), MoE (Rau, 2019), soft-mixture-of-experts (Puigcerver et al., 2023) and Transformers  (Wolf et al., 2020) library to implement our proposed GUSD.
-We conduct our experiments on a cluster with 4 Tesla V100 GPUs with 32 GB memory, 16 CPU cores, and 377GB CPU memory.
-Code, data, hyperparameters and trained models will be made publicly available.
-Appendix B Hyperparameters
+Table 5: Hyperparameter settings of GUSD in Kaggle
+Hyperparameter Kaggle LCS
+optimizer AdamW AdamW
+learning rate 1e-4 1e-4
+weight decay 1e-5 1e-5
+CrossEntropyLoss weight 1.3 1.3
+epochs 20 20
+dropout 0.3 0.3
+batch size 128 128
+early stop patience 3 3
+LM Bge-Large [41] Roberta [19]
+GNN layer 2 3
+GNN input dim 1024 768
+GNN hidden dim 1024 768
+GNN output dim 1024 768
+RetGAT hop num 2 2
+RetGAT hop decay $\alpha$ 0.3 0.3
+meta dim 3 6
+meta encoder hidden dim 1024 768
+meta encoder output dim 1024 768
+TRM heads 4 4
+TRM layers 2 2
+GMoE layer 2 2
+GMoE hidden dim 1024 768
+GMoE output dim 1024 768
+GMoE experts number 21 28
+Appendix 0.C Hyperparameters
 We present our hyperparameter settings in Table 5 to facilitate reproduction. Code, data, and trained models will be made publicly available.
-Appendix C k-hop Neighbors Algorithm
-Previous studies (Atwood and Towsley, 2016; Wang et al., 2020) have used the $k$-th power of the adjacency matrix to capture $k$-hop visibility.
-This method does not strictly define $k$-hop neighbors as nodes reachable in exactly k steps.
-To address this, we introduce a new GPU-friendly $k$-hop Algorithm that quickly and accurately computes $k$-hop neighbors using adjacency matrices.
-The details are illustrated in Algorithm 1.
-Considering that the number of $k$-hop neighbors $\mathcal{A}_{k}$ ($k>1$) can be large and may require substantial computational resources, we introduce an $r$-ratio (a hyper-parameter) to randomly sample neighbors from the entire set of neighbors for each hop.
 Algorithm 1 $k$-hop Neighbors with Minimum Distance
@@
 14:return $\mathcal{A}_{k}$
-Table 5. Hyperparameter settings of GUSD in Kaggle
-Hyperparameter Kaggle LCS
-optimizer AdamW AdamW
-learning rate 1e-4 1e-4
-weight decay 1e-5 1e-5
-CrossEntropyLoss weight 1.3 1.3
-epochs 20 20
-dropout 0.3 0.3
-batch size 128 128
-early stop patience 3 3
-LM Bge-Large (Xiao et al., 2023) Roberta (Liu, 2019)
-GNN layer 2 3
-GNN input dim 1024 768
-GNN hidden dim 1024 768
-GNN output dim 1024 768
-RetGAT hop num 2 2
-RetGAT hop decay $\alpha$ 0.3 0.3
-meta dim 3 6
-meta encoder hidden dim 1024 768
-meta encoder output dim 1024 768
-TRM heads 4 4
-TRM layers 2 2
-GMoE layer 2 2
-GMoE hidden dim 1024 768
-GMoE output dim 1024 768
-GMoE experts number 21 28
+Appendix 0.D k-hop Neighbors Algorithm
+Previous studies [1, 37] have used the $k$-th power of the adjacency matrix to capture $k$-hop visibility.
+This method does not strictly define $k$-hop neighbors as nodes reachable in exactly k steps.
+To address this, we introduce a new GPU-friendly $k$-hop Algorithm that quickly and accurately computes $k$-hop neighbors using adjacency matrices.
+The details are illustrated in Algorithm 1.
+Considering that the number of $k$-hop neighbors $\mathcal{A}_{k}$ ($k>1$) can be large and may require substantial computational resources, we introduce an $r$-ratio (a hyper-parameter) to randomly sample neighbors from the entire set of neighbors for each hop.
+Appendix 0.E Metadata Collected for Nodes in Datasets
+Here we provide the details of metadata collected for nodes in two datasets in Table 6
+Table 6: Comparison of metadata fields used in the LCS and IMDB datasets.
+Entity Field Type LCS IMDB
+Movie year int
+isAdult bool
+runtime int
+rating float
+numVotes int
+genres list
+User badge_count int
+review_count int
+avg_helpful float
+avg_total float
+avg_score float
+avg_rating float
+avg_length float
+Review helpful_votes int
+total_votes int
+score float
+rating float
+review_text_length int